if 1 is add to both numerator and and denominator of a fraction the fraction becomes 1/2. if 8 is added to both, the fraction become 2/3. what is d fraction

Answer by Alan3354(69443)

if 1 is add to both numerator and and denominator of a fraction the fraction becomes 1/2. if 8 is added to both, the fraction become 2/3. what is d fraction
-------------
(n+1)/(d+1) = 1/2
(n+8)/(d+8) = 2/3
---
2n + 2 = d + 1
3n + 24 = 2d + 16
4n + 4 = 2d + 2 ------ 1st eqn times 2
------------------------------------------------ Subtract
-n + 20 = 14
n = 6
d = 13